# Accelerating Focal Search in Multi-Agent Path Finding with Tighter Lower Bounds

### v1

@@ Abstract
Multi-Agent Path Finding (MAPF) involves finding collision-free paths for multiple agents while minimizing a cost function—an NP-hard problem. Bounded suboptimal methods like Enhanced Conflict-Based Search (ECBS) and Explicit Estimation CBS (EECBS) balance solution quality with computational efficiency using focal search mechanisms. While effective, traditional focal search faces a limitation: the lower bound (LB) value determining which nodes enter the FOCAL list often increases slowly in early search stages, resulting in a constrained search space that delays finding valid solutions. In this paper, we propose a novel bounded suboptimal algorithm, double-ECBS (DECBS), to address this issue by first determining the maximum LB value and then employing a best-first search guided by this LB to find a collision-free path. Experimental results demonstrate that DECBS outperforms ECBS in most test cases and is compatible with existing optimization techniques. DECBS can reduce nearly 30% high-level CT nodes and 50% low-level focal search nodes. When agent density is moderate to high, DECBS achieves a 23.5% average runtime improvement over ECBS with identical suboptimality bounds and optimizations.
@@ I INTRODUCTION
Solving the MAPF problem optimally is known to be NP-hard [7], creating significant challenges for optimal solvers regarding scalability and efficiency. In contrast, suboptimal MAPF solvers—including Prioritized Planning (PP) [8, 9], PBS [10], and their variants [11, 12]—offer improved computational performance but lack the theoretical guarantees on solution quality. Bounded-suboptimal approaches like Enhanced CBS (ECBS) [13] and EECBS [14] provide a trade-off between efficiency and solution quality, guaranteeing collision-free solutions’ costs are within a user-defined suboptimality factor of the optimal cost.
@@
The key idea behind bounded-suboptimal MAPF solvers lies in relaxing the best-first search requirement while incorporating additional heuristic functions to select the next node for expansion effectively. ECBS [13] achieves bounded suboptimality by replacing CBS’s best-first search strategy at both the high and low levels of CBS with focal search [15, 16]. Focal search operates through two queues: an OPEN list containing all unexpanded nodes sorted by cost, a FOCAL list containing a subset of these nodes. The algorithm selects expansion candidates from the FOCAL list using an auxiliary heuristic to accelerate finding bounded-suboptimal solutions. In our implementation, we use collision minimization as the FOCAL heuristic function at both search levels, explicitly employing the negative count of collisions in paths or complete solutions.
During focal search, since the optimal cost is unknown until an optimal solution is found, focal search uses the lowest cost in the OPEN list as a lower bound (LB) value to estimate the optimal cost and define FOCAL list eligibility. However, when the LB value is small, the FOCAL list contains fewer nodes, causing focal search to behave similarly to standard best-first search. This constraint also leads to unbalanced search tree expansion [14].
@@
In this paper, we introduce double-ECBS (DECBS), a novel bounded-suboptimal variant of ECBS. DECBS modifies the original focal search used in low-level path planning by incorporating an optimized search manner, what we call double search. This approach is inspired by ITA-ECBS [17], a bounded-suboptimal version of ITA-CBS [18]. However, ITA-ECBS primarily focuses on integrating ECBS and ITA-CBS, and it has not been developed as a universal optimization applicable to other methods. The double search operates in two phases: first, a shortest path search determines the optimal cost, the upper bound for the LB value; second, DECBS employs a best-first search guided by the FOCAL list auxiliary heuristic function using this LB value to find a path with fewer collisions. Importantly, paths returned during the low-level search are not necessarily collision-free—they need only to satisfy the constraints specified in the corresponding high-level CT node.
@@
This maximum LB value offers three benefits: (1) by accurately reflecting the optimal cost, it enables more nodes to enter the FOCAL list. This increases the likelihood of selecting a path with fewer collisions, thereby reducing both high-level node expansion and low-level node expansion; (2) The node of focal search is more expensive than $A^{*}$ because it requires computing an additional heuristic. In double search, we first perform an $A^{*}$ search to obtain a larger LB. This larger LB, in turn, reduces the number of nodes generated by the focal search. Although the $A^{*}$ search introduces additional nodes, these nodes are cheaper to compute than the focal search nodes. (3) Candidate nodes with costs exceeding the suboptimality can be pruned, saving memory during search operations. In summary, although DECBS performs two path searches for each low-level search, the double search is faster than executing a single focal search. Our empirical evaluation confirms these benefits, showing that DECBS reduces nearly 30% high-level CT nodes and 50% low-level focal search nodes, and achieves a 23.5% average runtime improvement over ECBS with identical optimizations when agent density is moderate to high.
@@ III-A Focal Search
The OPEN list stores all candidate nodes awaiting expansion, sorted by the function $f(n)=g(n)+h(n)$, where $g(n)$ is the cost-so-far and $h(n)$ is an admissible heuristic estimate of cost-to-go, similar to $A^{*}$ search. The FOCAL queue contains only those nodes $n$ from OPEN satisfying $f(n)\leq w\cdot f_{front}$, where $f_{front}$ is the minimum $f$ value in OPEN. Nodes in FOCAL are sorted according to an additional heuristic function $d(n)$ that guides the search toward promising regions of the search space.
Focal search always expands the front node in the FOCAL list. When a solution with cost $c^{val}$ is found, the algorithm sets the lower bound (LB) $c^{g}=f_{front}$. We can observe that $f_{front}\leq c^{opt}$ and $f_{front}\leq c^{val}\leq wf_{front}\leq wc^{opt}$. Finally, focal search returns two outputs: the LB $c^{g}$ and the solution with cost $c^{val}$. If no solution exists, both $c^{g}$ and $c^{val}$ are set to $\infty$.
@@ III-B Multi-Agent Path Finding (MAPF)
Multi-Agent Path Finding (MAPF) has a long history [9]. Many algorithms have been developed to solve it or its variants. Decoupled algorithms [9, 19, 20] plan paths for each agent independently and combine these paths into a single solution. Coupled algorithms [21, 22] plan simultaneously for all agents. There also exist dynamically-coupled algorithms [4, 23] initially generate individual paths and re-plan for multiple agents together when resolving collisions. Among these approaches, Conflict-Based Search (CBS) [4] is a basic centralized optimal MAPF algorithm. Some bounded-suboptimal algorithms, such as ECBS [13] and EECBS [14], are based on it.
@@
Conflict-Based Search (CBS) is an optimal two-level search algorithm. The low-level computes shortest paths for individual agents from their start locations to goals, while the high-level searches a binary Constraint Tree (CT). Each CT node $H=(c,\Omega,\pi)$ includes a constraint set $\Omega$, a solution $\pi$ containing shortest paths satisfying $\Omega$ for all agents, and the solution cost $c$. When a solution $\pi$ or a path does not include any agent actions or positions that are restricted by a $\Omega$, we say this solution or path satisfies the $\Omega$. Importantly, a solution $\pi$ satisfying $\Omega$ may still contain collisions; we consider $\pi$ a valid solution pnly when it is collision-free.
@@
BCBS [13] and ECBS [13] build upon CBS by incorporating focal search. In ECBS low-level search, focal search returns both an LB value $c^{g}_{i}$ and a valid path $\pi_{i}$, with cost $c_{i}$ for agent $i$, satisfying: $c^{g}_{i}\leq c_{i}^{opt}\leq c_{i}\leq w\cdot c^{g}_{i}$, where $c_{i}^{opt}$ represents the optimal path cost of agent $i$ under constraints $\Omega$. At the high-level search, ECBS extends the CT node structure to $H=(c,\Omega,\pi,L,c_{L})$, adding an array $L$, that stores all LB values $c^{g}_{i}$ for paths in $\pi$, and cost $c_{L}$ represents the sum of these lower bounds. ECBS’s high-level search maintains two priority queues: FOCAL and OPEN. OPEN sorts all CT nodes by ascending $c_{L}$, while FOCAL contains nodes $H$ from OPEN satisfying $H.c\leq w\cdot H_{front}.c_{L}$, where $H_{front}$ is the front CT node in OPEN. FOCAL is ordered by a user-defined heuristic function $d(H)$. ECBS guarantees that the returned solution $H^{sol}.\pi$ satisfies $H^{sol}.c\leq w\cdot c^{opt}$, with $c^{opt}$ being the optimal valid solution cost. BCBS similarly employs focal search at both levels but defines the high-level CT node LB value as the sum of the low-level path costs. BCBS is typically denoted as BCBS$(w_{1},w_{2})$, where $w_{1}$ and $w_{2}$ are the suboptimality factors for high- and low-level search. When $w_{1}w_{2}=w$, BCBS$(w_{1},w_{2})$ returns a solution that is bounded-suboptimal with factor $w$. Notably, BCBS$(w,1)$ is equivalent to using $A^{*}$ in low-level and focal search in high-level.
@@ IV Method
During focal search, nodes from OPEN with costs in the range $[f_{front},wf_{front}]$ are selected into the FOCAL. If we want to include more nodes in FOCAL (e.t. potentially find solutions more quickly), we must increase $f_{front}$ by expanding the front node in OPEN. However, this process is often inefficient because $f_{front}$ increases slowly–it only changes after exhausting all existing and potential child nodes with the current minimum cost.
We observe that a more effective range would be $[f_{front},wc^{opt}]$, as indicated by the inequality. Since $c^{opt}$ represents as an upper bound for $f_{front}$, this range maxmimizes the number of candidates for FOCAL selection while still maintaining the theoretical suboptimality guarantee. The challenge of using $[f_{front},wc^{opt}]$ is that $c^{opt}$ remains unknown during focal search. Determine $c^{opt}$ requires running an optimal path finding algorithm like $A^{*}$. So in our double search approach, we first run shortest path search such as $A^{*}$ to obtain $c^{opt}$. Once we have $c^{opt}$, we no longer need OPEN from focal search, as we can immediately discard any node with cost exceeding $wc^{opt}$, since they won’t in a bound-suboptimal solution. This allows us to conduct a best-first search exclusively on the FOCAL, using conflict number as the heuristic function to sort all CT nodes.
Algorithm 1 presents the pseudo-code for DECBS. The key component appears in Lines 23–26: first, we determine the lower bound $c^{g}$ through a shortest path search, then use this value to filter out nodes exceeding the suboptimality bound before adding the remaining nodes into the FOCAL. Finally, we expand nodes from FOCAL according to the the FOCAL heuristic function.
@@
Input: Graph $G$, start locations $\{s_{i}\}$, target locations $\{g_{i}\}$, target matrix $A$, suboptimality factor $w$
@@
16:             $Q.\Omega$ = $Q.\Omega$cup ($k$, $v^{k}_{t}$, $t$) // vertex constraint
@@
18:             $Q.\Omega$ = $Q.\Omega$cup ($k$, $v^{k}_{t-1}$, $v^{k}_{t}$, $t$) // edge constraint          $Q.\pi,Q.L$ = lowLevelSearch($G,s_{k},g_{x},Q,w$)
@@
People may have two questions, since running an optimal search followed by a bounded-suboptimal search seems wasteful and inefficient–
  1. 1.
    Why perform two searches when the first already yields a valid path?
  2. 2.
    Does double search more efficient than single focal search?
@@
For the second question, although performing a double search might seem slower than a single search, three factors make our approach more efficient: (1) By obtaining $c^{opt}$, we can include more nodes in the FOCAL list compared to ECBS. A larger FOCAL set increases the likelihood of finding a bounded-suboptimal path with fewer conflicts, thereby reducing the number of nodes expanded during the low-level search. We show an idea example in Figure 1 to explain this situation. (2) Since we already have an upper bound for each agent’s path LB value, the CT node LB becomes larger. This larger LB provides the high-level focal search with greater flexibility in selecting FOCAL nodes for expansion, resulting in fewer CT nodes overall. (3) Focal search is more computationally expensive than $A^{*}$ because it requires evaluating an additional heuristic. The larger LB reduces the number of nodes generated by the focal search. Although the $A^{*}$ search introduces extra nodes, these nodes are cheaper to compute than those produced by the focal search. In effect, the additional $A^{*}$ search helps offset the cost of the extra nodes, leading to an overall reduction in computation time.
Figure 2: DECBS and ECBS Comparison of Search Nodes: Only test cases where both algorithms finished within the time limit were counted. The top figure shows the number of low-level search nodes—focal search expanded nodes for ECBS and best-first search nodes for DECBS. The bottom figure displays the number of expanded CT nodes.
@@ V Experimental Results
In our experiments, we evaluate DECBS against ECBS, including variants with two optimization techniques, Bypassing Conflicts (BC) [24] and Target Reasoning (TR) [25]. We implement DECBS and ECBS in Rust.222 Our code is already available at https://github.com/HarukiMoriarty/RUST-CBS. All experiments are conducted on an Ubuntu 20.04.1 system with an AMD Ryzen 3990X 64-Core Processor with 2133 MHz 196GB RAM.
Figure 4: In this figure, we show average DECBS runtime improvements over ECBS on the same test cases and using the same optimization methods. We calculate improvement as follows: $(runtime(ECBS)-runtime(DECBS))/runtime(ECBS)$. Columns 1 and 3 represent improvements for different agent numbers, while columns 2 and 4 represent improvements for different suboptimality factors. We only selected test cases that were solved by both algorithms.
@@ V-A Test Settings
We evaluated DECBS and ECBS on 6 maps from the MAPF Benchmark [26], as shown in Figure 3: (1) random-32-32-20 (32x32) and maze-32-32-20 (32x32) are grid maps with random obstacles and designed obstacles, (2) den_312d (65x81) is a grid map from the video game Dragon Age Origins, (3) warehouse-10-20-10-2-1 (161x63) is a grid map inspired by real-world autonomous warehouses, and (4) den_520d (256x257) and Paris_1_256 (256x256) are large benchmark maps. We use map-specific suboptimality factors: random-32-32-20 and empty-32-32-20 (1.02 to 1.20, interval 0.02), den_312d and warehouse-10-20-10-2-1 (1.01 to 1.10, interval 0.01), den_520d and Paris_1_256 (1.002 to 1.038, interval 0.004). For agent numbers, we have empty-32-32-20 (10 to 80, interval 10), random-32-32-20 (45 to 150, interval 15), den_312d (105 to 210, interval 15), warehouse-10-20-10-2-1 (90 to 300, interval 30), den_520d and Paris_1_256 (100 to 450, interval 50).
For each combination of map, agent number, and suboptimality factor, we generate 200 random test cases, resulting in a total of 270,000 test cases. An algorithm is considered to have failed for a given test case if it does not find a valid solution within 60 seconds (with runtime recorded as 60 seconds). Success rate is defined as the percentage of test cases solved successfully within the time limit under specific test settings.

### v2

@@ Abstract
Multi-Agent Path Finding (MAPF) involves finding collision-free paths for multiple agents while minimizing a cost function—an NP-hard problem. Bounded suboptimal methods like Enhanced Conflict-Based Search (ECBS) and Explicit Estimation CBS (EECBS) balance solution quality with computational efficiency using focal search mechanisms. While effective, traditional focal search faces a limitation: the lower bound (LB) value determining which nodes enter the FOCAL list often increases slowly in early search stages, resulting in a constrained search space that delays finding valid solutions. In this paper, we propose a novel bounded suboptimal algorithm, double-ECBS (DECBS), to address this issue by first determining the maximum LB value and then employing a best-first search guided by this LB to find a collision-free path. Experimental results demonstrate that DECBS outperforms ECBS in most test cases and is compatible with existing optimization techniques. DECBS can reduce nearly 30% high-level Constraint Tree (CT) nodes and 50% low-level focal search nodes. When agent density is high, DECBS achieves a 23.5% average runtime improvement over ECBS with identical suboptimality bounds and optimizations.
@@ I INTRODUCTION
Solving the MAPF problem optimally is known to be NP-hard [7], creating significant challenges for optimal solvers regarding scalability and efficiency. In contrast, suboptimal MAPF solvers—including Prioritized Planning (PP) [8, 9], PBS [10], and their variants [11, 12]—offer improved computational performance but lack the theoretical guarantees on solution quality. There also exists asymptotically optimal method such as LaCAM [13]. Bounded-suboptimal approaches like Enhanced CBS (ECBS) [14] and EECBS [15] provide a trade-off between efficiency and solution quality, guaranteeing collision-free solutions’ costs are within a user-defined suboptimality factor of the optimal cost.
@@
The key idea behind bounded-suboptimal MAPF solvers lies in relaxing the best-first search requirement while incorporating additional heuristic functions to select the next node for expansion effectively. ECBS [14] achieves bounded suboptimality by replacing CBS’s best-first search strategy at both the high and low levels of CBS with focal search [16, 17]. Focal search operates through two queues: an OPEN list containing all unexpanded nodes sorted by cost, a FOCAL list containing a subset of these nodes. The algorithm selects expansion candidates from the FOCAL list using an auxiliary heuristic to accelerate finding bounded-suboptimal solutions. In our implementation, we use collision minimization as the FOCAL heuristic function at both search levels, explicitly employing the negative count of collisions in paths or complete solutions. During focal search, since the optimal cost is unknown until an optimal solution is found, focal search uses the lowest cost in the OPEN list as a lower bound (LB) value to estimate the optimal cost and define FOCAL list eligibility. However, when the LB value is small, the FOCAL list contains fewer nodes, causing focal search to behave similarly to standard best-first search. This constraint also leads to unbalanced search tree expansion [15].
@@
In this paper, we introduce double-ECBS (DECBS), a novel bounded-suboptimal variant of ECBS. DECBS modifies the original focal search used in low-level path planning by incorporating an optimized search manner, what we call double search. This approach is inspired by ITA-ECBS [18], a bounded-suboptimal version of ITA-CBS [19]. However, ITA-ECBS primarily focuses on integrating ECBS and ITA-CBS, and it has not been developed as a universal optimization applicable to other methods. The double search operates in two phases: first, a shortest path search determines the optimal cost, the upper bound for the LB value; second, DECBS employs a best-first search guided by the FOCAL list auxiliary heuristic function using this LB value to find a path with fewer collisions. Importantly, paths returned during the low-level search are not necessarily collision-free—they need only to satisfy the constraints specified in the corresponding high-level Constraint Tree (CT) node.
@@
This maximum LB value offers three benefits: (1) by accurately reflecting the optimal cost, it enables more nodes to enter FOCAL list. This increases the likelihood of selecting a path with fewer collisions, thereby reducing both high-level node expansion and low-level node expansion; (2) The node of focal search is more expensive than $A^{*}$ because it requires computing an additional heuristic. In double search, we first perform an $A^{*}$ search to obtain a larger LB. This larger LB, in turn, reduces the number of nodes generated by focal search. Although $A^{*}$ search introduces additional nodes, these nodes are cheaper to compute than focal search nodes. (3) Candidate nodes with costs exceeding the suboptimality can be pruned, saving memory during search operations. In summary, although DECBS performs two path searches for each low-level search, double search is faster than executing a single focal search. Our empirical evaluation confirms these benefits, showing DECBS reduces nearly 30% high-level CT nodes and 50% low-level focal search nodes, and achieves a 23.5% average runtime improvement over ECBS with identical optimizations when agent density is high.
@@ III-A Focal Search
The OPEN list stores all candidate nodes awaiting expansion, sorted by the function $f(n)=g(n)+h(n)$, where $g(n)$ is the cost-so-far and $h(n)$ is an admissible heuristic estimate of cost-to-go, similar to $A^{*}$ search. The FOCAL queue contains only those nodes $n$ from OPEN satisfying $f(n)\leq w\cdot f_{front}$, where $f_{front}$ is the minimum $f$ value in OPEN. Nodes in FOCAL are sorted according to an additional heuristic function $d(n)$ that guides the search toward promising regions of the search space. Focal search always expands the front node in the FOCAL list. When a solution with cost $c^{val}$ is found, the algorithm sets the lower bound (LB) $c^{g}=f_{front}$. We can observe that $f_{front}\leq c^{opt}$ and $f_{front}\leq c^{val}\leq wf_{front}\leq wc^{opt}$. Finally, focal search returns two outputs: the LB $c^{g}$ and the solution with cost $c^{val}$. If no solution exists, both $c^{g}$ and $c^{val}$ are set to $\infty$.
@@ III-B Multi-Agent Path Finding (MAPF)
Multi-Agent Path Finding (MAPF) has a long history [9]. Many algorithms have been developed to solve it or its variants. Decoupled algorithms [9, 20, 21] plan paths for each agent independently and combine these paths into a single solution. Coupled algorithms [22, 23] plan simultaneously for all agents. There also exist dynamically-coupled algorithms [4, 24, 13] initially generate individual paths and re-plan for multiple agents together when resolving collisions. Among these approaches, Conflict-Based Search (CBS) [4] is a basic centralized optimal MAPF algorithm. Some bounded-suboptimal algorithms, such as ECBS [14] and EECBS [15], are based on it.
@@
Conflict-Based Search (CBS) is an optimal two-level search algorithm. The low-level computes shortest paths for individual agents from their start locations to goals, while the high-level searches a binary Constraint Tree. Each CT node $H=(c,\Omega,\pi)$ includes a constraint set $\Omega$, a solution $\pi$ containing shortest paths satisfying $\Omega$ for all agents, and the solution cost $c$. When a solution $\pi$ or a path does not include any agent actions or positions that are restricted by a $\Omega$, we say this solution or path satisfies the $\Omega$. Importantly, a solution $\pi$ satisfying $\Omega$ may still contain collisions; we consider $\pi$ a valid solution only when it is collision-free.
@@
[…]
@@ IV Method
During focal search, nodes from OPEN with costs in the range $f_{front},wf_{front}]$ are selected into the FOCAL. If we want to include more nodes in FOCAL (e.t. potentially find solutions more quickly), we must increase $f_{front}$ by expanding the front node in OPEN. However, this process is often inefficient because $f_{front}$ increases slowly–it only changes after exhausting all existing and potential child nodes with the current minimum cost.
We observe that a more effective range would be $f_{front},w·c^{opt}]$, as indicated by the inequality. Since $c^{opt}$ represents an upper bound for $f_{front}$, this range maximizes the number of candidates for FOCAL selection while still maintaining the theoretical suboptimality guarantee. The challenge of using $f_{front},wc^{opt}]$ is that $c^{opt}$ remains unknown during focal search. Determining $c^{opt}$ requires running an optimal path finding algorithm like $A^{*}$. So in our double search approach, we first run shortest path search such as $A^{*}$ to obtain $c^{opt}$. Once we have $c^{opt}$, we no longer need OPEN from focal search, as we can immediately discard any node with cost exceeding $wc^{opt}$, since they won’t be in a bound-suboptimal solution. This allows us to conduct a best-first search exclusively on the FOCAL, using conflict number as the heuristic function to sort all CT nodes.
Algorithm˜1 presents the pseudo-code for DECBS. The key component appears in Lines 23–26: first, we determine the lower bound $c^{g}$ through a shortest path search, then use this value to filter out nodes exceeding the suboptimality bound before adding the remaining nodes into the FOCAL. Finally, we expand nodes from FOCAL according to the the FOCAL heuristic function.
@@
Input: Graph $G$, start locations $\{s_{i}\}$, target locations $\{g_{i}\}$, suboptimality factor $w$
@@
16:     $Q.\Omega$ = $Q.\Omega$ $\cup$ ($k$, $v^{k}_{t}$, $t$) // vertex constraint
@@
18:     $Q.\Omega$ = $Q.\Omega$ $\cup$ ($k$, $v^{k}_{t-1}$, $v^{k}_{t}$, $t$) // edge constraint     $Q.\pi,Q.L$ = lowLevelSearch($G,s_{k},g_{x},Q,w$)
@@
People may have two questions, since running an optimal search followed by a bounded-suboptimal search seems wasteful and inefficient–(1) Why perform two searches when the first already yields a valid path? (2) Is double search more efficient than single focal search?
@@
For the second question, although performing a double search might seem slower than a single search, three factors make our approach more efficient: (1) By obtaining $c^{opt}$, we can include more nodes in the FOCAL list compared to ECBS. A larger FOCAL set increases the likelihood of finding a bounded-suboptimal path with fewer conflicts, thereby reducing the number of nodes expanded during the low-level search. We show an idea example in Figure˜1 to explain this situation. (2) Since we already have an upper bound for each agent’s path LB value, the CT node LB becomes larger. This larger LB provides the high-level focal search with greater flexibility in selecting FOCAL nodes for expansion, resulting in fewer CT nodes overall. (3) Focal search is more computationally expensive than $A^{*}$ because it requires evaluating an additional heuristic. The larger LB reduces the number of nodes generated by the focal search. Although $A^{*}$ search introduces extra nodes, these nodes are cheaper to compute than those produced by focal search. In effect, the additional $A^{*}$ search helps offset the cost of extra nodes, leading to an overall reduction in computation time.
Figure 2: Comparison of Search Nodes between DECBS and ECBS: Points above the blue line (1x) indicate that DECBS expands fewer search nodes than ECBS. Only test cases where both algorithms finished within the time limit were counted. The top figure shows the number of low-level search nodes—focal search expanded nodes for ECBS and best-first search nodes for DECBS. The bottom figure displays the number of expanded CT nodes.
@@ V Experimental Results
In our experiments, we evaluate DECBS against ECBS, including variants with two optimization techniques, Bypassing Conflicts (BC) [25] and Target Reasoning (TR) [26]. We implement DECBS and ECBS in Rust.222 Our code is already available at https://github.com/HarukiMoriarty/RUST-CBS. To ensure a fair comparison, both DECBS and ECBS share the same code base for core components such as low-level focal search and high-level search node expansion. All experiments are conducted on an Ubuntu 20.04.1 system with an AMD Ryzen 3990X 64-Core Processor with 2133 MHz 196GB RAM.
Figure 4: In this figure, we show average DECBS runtime improvements over ECBS on the same test cases and using the same optimization methods. We calculate improvement as follows: (runtime(ECBS)runtime(DECBS))/runtime(ECBS)(runtime(ECBS)-runtime(DECBS))/runtime(ECBS). Columns 1 and 3 represent improvements for different agent numbers, while columns 2 and 4 represent improvements for different suboptimality factors. We only selected test cases that were solved by both algorithms.
@@ V-A Test Settings
We evaluated DECBS and ECBS on 6 maps from the MAPF Benchmark [27], as shown in Figure˜3: (1) random-32-32-20 (32x32) and maze-32-32-20 (32x32) are grid maps with random obstacles and designed obstacles, (2) den_312d (65x81) is a grid map from the video game Dragon Age Origins, (3) warehouse-10-20-10-2-1 (161x63) is a grid map inspired by real-world autonomous warehouses, and (4) den_520d (256x257) and Paris_1_256 (256x256) are large benchmark maps. We use map-specific suboptimality factors: random-32-32-20 and empty-32-32-20 (1.02 to 1.20, interval 0.02), den_312d and warehouse-10-20-10-2-1 (1.01 to 1.10, interval 0.01), den_520d and Paris_1_256 (1.002 to 1.038, interval 0.004). For agent numbers, we have empty-32-32-20 (10 to 80, interval 10), random-32-32-20 (45 to 150, interval 15), den_312d (105 to 210, interval 15), warehouse-10-20-10-2-1 (90 to 300, interval 30), den_520d and Paris_1_256 (100 to 450, interval 50). For each combination of map, agent number and suboptimality factor, we generate 200 random test cases, resulting in total 270,000 test cases. An algorithm is considered to have failed for a given test case if it does not find a valid solution within 60 seconds (with runtime recorded as 60 seconds). Success rate is defined as the percentage of test cases solved successfully within the time limit under specific test settings.